# Improving LLM Unlearning Robustness via Random Perturbations

### v3

@@ Abstract
In this paper, we show that current state-of-the-art LLM unlearning methods inherently reduce models’ robustness, causing them to misbehave even when a single non-adversarial forget-token is in the retain-query. Toward understanding underlying causes, we reframe the unlearning process as backdoor attacks and defenses: forget-tokens act as backdoor triggers that, when activated in retain-queries, cause disruptions in unlearned models’ behaviors, similar to successful backdoor attacks. To mitigate this vulnerability, we propose Random Noise Augmentation (RNA)—a plug-and-play, model and method agnostic approach with theoretical guarantees for improving the robustness of unlearned models. Extensive experiments demonstrate that RNA significantly improves the robustness of unlearned models, maintains unlearning performances while introducing no additional computational overhead.
@@ 1 Introduction
Modern LLMs are pre-trained on massive text corpora and then post-trained with Reinforcement Learning from Human Feedback (RLHF; [9, 68, 54, 41]) to be helpful and harmless [2]. Recent studies have shown that despite safety enhancements, aligned LLMs can still exhibit harmful and undesirable behaviors, such as generating toxic content [62], producing copyrighted material [26, 14, 61], gender bias [4], leaking sensitive and private information [39, 43], and potentially aiding malicious uses such as cyberattacks, chemical attacks, and bioweapons development [17, 49, 30]. As LLMs advance in size and capabilities at an unprecedented speed, concerns about their potential risks continue to grow.
Machine Unlearning (MU; [7, 5, 40, 63, 10]) is an approach aiming to robustly (1) remove or suppress specific target knowledge in a forget-set and capabilities from a pre-trained model, while (2) retaining the model’s other knowledge in a retain-set and capabilities. Recent works on the robustness of unlearning methods primarily focus on the first criterion, evaluating the robustness of unlearned models against knowledge recovery that adversarially tries to recover unlearned knowledge. For example, previously unlearned knowledge is shown to resurface through relearning [30, 12, 34, 32], sequential unlearning [52], target relearning attacks [21], removing/steering specific directions in the latent space [34, 50], quantization [67], or even simply fine-tuning on unrelated tasks [13, 34].
However, the equally important criterion of robustly preserving the model’s general knowledge—that is, ensuring stable and accurate responses to retain-queries even when they inadvertently include forget-tokens—remains underexplored. Initial steps have been taken, such as Thaker et al. [57], who examined the robustness of Representation Misdirection for Unlearning (RMU; [30]), demonstrating that RMU-unlearned models are fragile when asked with retain-queries (e.g., Q&A about general knowledge) containing forget-tokens (tokens in the forget-set). However, many critical questions remain unanswered. In this paper, we make the following contributions:
Figure 1: Top-left: Accuracy of RM and RM w/ RNA models on MMLU [20] and perturbed MMLU (MMLU QA contains forget-tokens; see Appendix A for details). Top-right: Accuracy of PO and PO w/ RNA models on MMLU and perturbed MMLU. Bottom-left: Accuracy of all unlearned models on WMDP [30] and perturbed WMDP. Bottom-right: Accuracy of all unlearned models on MMLU College Biology and Computer Security. Original RM models are shown by one-color circles and original PO models by one-color triangles. Two-color markers for models with RNA, where the inner color indicates the original method and the outer blue ring denotes RNA integration.
@@
➁ We propose a novel perspective that decomposes the unlearning process into “forgetting” and “retaining” processes and reframes it as a backdoor attack and defense problem. The “forgetting” corresponds to a backdoor attack: by treating the forget-set as a poisoned dataset, we formulate how the unlearning methods inadvertently learn to align forget-tokens (backdoor triggers) with the target representations (the target labels). As a result, when forget-tokens appear in a retain-query, it is similar to activating the backdoor trigger making the model misbehave. To counteract the vulnerability introduced by the “forgetting”, we reinterpret the “retaining” as a backdoor defense and introduce Random Noise Augmentation (RNA), a plug-and-play, model- and method-agnostic approach which adds small, independent Gaussian noise to each retain-query’s representation during fine-tuning to reduce the model’s sensitivity to forget-tokens.
➂ Through theoretical and empirical analysis, we show that RNA significantly improves the robustness of unlearned models while maintaining original forget and retain performances (Figure 1).
@@ 2.1 Related Works
Machine Unlearning. MU has become one of the most important tools for ensuring the safety and protecting the privacy of LLMs [63, 40]. Most recent works on MU focus on developing algorithms for different tasks, domains, and settings [44, 58, 25, 52, 8, 42, 38, 60, 28], while much less effort was spent on developing robust unlearning algorithms.
@@
Unlearning Robustness. Previous works on MU robustness focus on “forget-robustness”, studying the robustness of MU algorithms in making the model forget the target knowledge and capabilities. Researchers showed that unlearned knowledge can resurface through re-learning [30, 35, 3, 32], sequential unlearning [52], quantization [67], fine-tuning unlearned models on unrelated tasks [13, 34], and adversarial attacks [21, 64, 53, 22] and developed methods for improving forget-robustness of MU algorithms [51, 56, 55, 16, 67]. This paper explores the “retain-robustness” of MU algorithms, studying the robustness of MU algorithms in robustly retaining the original model’s general knowledge and capabilities. Thaker et al. [57] presented preliminary results showing that state-of-the-art MU algorithms do not preserve the original model’s knowledge and capabilities. We bridge the gap in retain-robustness research by introducing RNA, a simple data augmentation method inspired by adversarial training to improve the robustness of MU algorithms.
2.2 Preliminaries
@@
Notation and problem formulation. The training data of an MU problem consists of two subsets: the forget-set $\mathcal{D}_{f}$ and the retain-set $\mathcal{D}_{r}$. The goal is to minimize the model’s performance on the forget set while keeping the performance on the retain-set. Let $f_{\theta}$ be a model parameterized by $\theta$, and $\ell(y|x;\theta)$ is the loss of input $x$ with respect to a target output $y$ in model $f_{\theta}$. A commonly used form of unlearning involves minimizing the following two-part loss:
$\mathcal{L}_{\mathcal{D}_{f},\mathcal{D}_{r},\theta}=\mathbb{E}_{(x^{f},y^{f})\sim\mathcal{D}_{f}}\left[\ell\left(y^{f}|x^{f};\theta\right)\right]+\alpha\mathbb{E}_{(x^{r},y^{r})\sim\mathcal{D}_{r}}\left[\ell\left(y^{r}|x^{r};\theta\right)\right]$ (1)
@@
where $y^{f},\ y^{r}$ are the target outputs of forget and retain input, respectively, $\alpha\in\mathbb{R}^{+}$ is a retain weight. We consider two widely used classes of LLM unlearning methods, which rely on Representation Misdirection (RM) and Preference Optimization (PO).
2.2.1 Representation Misdirection
@@
Representation Misdirection (RMU and variants) are unlearning methods that achieve unlearning by manipulating latent representations during fine-tuning. Denote $\mathbf{z}_{\theta}^{f},\mathbf{z}_{\theta}^{r}\in\mathbb{R}^{n\times d_{l}}$ the output hidden state of $n$-tokens in forget-sample $x^{f}$ and in retain-sample $x^{r}$, respectively, at layer $l$ in model $f_{\theta}$, where $d_{l}$ is the dimension of hidden states at layer $l$. In this context, the activations from the MLP module in the transformer layer is used as the latent representation.
@@
Representation Misdirection for Unlearning (RMU; [30]) pushes the latent representation of forget-tokens to a predetermined random representation $y^{f}=c\mathbf{u}$, where $\mathbf{u}\in\mathbb{R}^{d_{l}}$ is a unit vector with each element uniformly sampled from $[0,1)$, and $c\in\mathbb{R}^{+}$ is a coefficient. It also regularizes the latent representation of retain-tokens back to the reference model’s representation:
$\mathcal{L}^{\text{RMU}}=\mathbb{E}_{x^{f}\sim\mathcal{D}_{f}}||\mathbf{z}_{\theta}^{f}-c\mathbf{u}||_{2}^{2}+\alpha\mathbb{E}_{x^{r}\sim\mathcal{D}_{r}}||\mathbf{z}_{\theta}^{r}-\mathbf{z}_{\theta^{\text{ref}}}^{r}||_{2}^{2},$ (2)
@@
where $\theta$ and $\theta^{\text{ref}}$ are the parameters of the updated and reference (frozen weight) models, respectively, and $\alpha\in\mathbb{R}^{+}$ is a retain weight.
Adaptive RMU  [11] is a variant of RMU that adaptively changes the coefficient of random vector $\mathbf{u}$ in the forget-loss based on the norm of forget-sample in the reference model. The target random representation $y^{f}=\beta||\mathbf{z}^{f}_{\theta^{\text{ref}}}||_{2}\mathbf{u}$, where $\beta\in\mathbb{R}^{+}$ is a scaling factor.
@@
Random Steering Vector (RSV). Additionally, we implement RSV—a variant of RMU that uses the target random representation $y^{f}=\mathbf{z}^{f}_{\theta^{\text{ref}}}+c\bm{\epsilon}$, where $c\in\mathbb{R}^{+}$ is a predetermined coefficient, $\bm{\epsilon}$ is a random unit vector sampled from Gaussian distribution $\mathcal{N}(\bm{0},\mu\bm{I})$, $\mu\bm{I}$ is covariance matrix, $\mu\in\mathbb{R}^{+}$.
2.2.2 Preference Optimization
@@
Negative Preference Optimization (NPO; [66]) treats forget-samples as negative preference samples in Direct Preference Optimization framework (DPO; [46]). NPO can be viewed as a gradient ascent variant with adaptive gradient weights that allows more controlled and stable optimization:
@@
$\mathcal{L}^{\text{NPO}}=\mathbb{E}_{(x^{f},y^{f})\sim\mathcal{D}_{f}}\left[-\frac{2}{\beta}\log\sigma\left(-\beta\log\left(\frac{\pi_{\theta}(y^{f}|x^{f})}{\pi_{\theta^{\text{ref}}}(y^{f}|x^{f})}\right)\right)\right],$ (3)
@@
where $\beta\in\mathbb{R}^{+}$ is a temperature hyperparameter (NPO reduces to gradient ascent as $\beta\to 0$), $\sigma(\cdot)$ is the sigmoid function, and $\pi_{\theta}(y^{f}|x^{f})$, $\pi_{\theta^{\text{ref}}}(y^{f}|x^{f})$ denotes the predicted probability of $y^{f}$ given $x^{f}$ in the model $f_{\theta}$ and reference model $f_{\theta^{\text{ref}}}$ (frozen weight) respectively.
@@
Simple Negative Preference Optimization (SimNPO; [15]) simplifies NPO by using a normalized sequence log-probability and introducing a reward margin hyperparameter $\gamma\geq 0$:
@@
$\mathcal{L}^{\text{SimNPO}}=\mathbb{E}_{(x^{f},y^{f})\sim\mathcal{D}_{f}}\left[-\frac{2}{\beta}\log\sigma\left(-\frac{\beta}{|y^{f}|}\log\pi_{\theta}(y^{f}|x^{f})-\gamma\right)\right],$ (4)
where $|y^{f}|$ is the length of output sequence $y^{f}$.
@@
Direct Preference Optimization (DPO; [46]). As a baseline, Zhang et al. [66], Maini et al. [36], Yuan et al. [65]) adopted standard DPO, using a refusal answer $y^{\text{idk}}\in\mathcal{D}_{\text{idk}}$ such as “I Don’t Know” as the positive samples and forget-samples as negative samples.
To preserve general capabilities, we use Mean Squared Error (MSE): $\mathcal{L}^{\text{MSE}}=\mathbb{E}_{(x^{r},y^{r})\sim\mathcal{D}_{r}}||\log\pi_{\theta}(x^{r})-\log\pi_{\theta^{\text{ref}}}(x^{r})||_{2}^{2}$ or Kullback–Leibler divergence (KL) $\mathcal{L}^{\text{KL}}=\mathbb{E}_{(x^{r},y^{r})\sim\mathcal{D}_{r}}\text{KL}\left(\log\pi_{\theta}(x^{r}),\log\pi_{\theta^{\text{ref}}}(x^{r})\right)$ as the retain-loss. Combining the two losses, we investigate a series of $6$ PO-based unlearning methods including NPO+MSE, NPO+KL, DPO+MSE, DPO+KL, SimNPO+MSE, and SimNPO+KL.
@@ 3 A Unified View of LLM Unlearning
We first draw a connection between RM and PO methods through a unified view of the generative latent variable model (GLVM). Let $\mathbf{z}_{\theta}^{f}+\mathbf{v}$ be the steered latent representation of forget-sample $x^{f}$ in $f_{\theta}$ as a result of RM. We assume that random vector $\mathbf{v}$ is small and sampled from normal distribution $\mathcal{N}(\bm{0},\mu\bm{I})$, $\mu\in\mathbb{R}^{+}$. We employ the notation of the GLVM, that is, GLVM $f_{\theta}$ generates target output $y^{f}$ given the latent variable $\mathbf{z}_{\theta}^{f}$. Let $\ell(y^{f}|\mathbf{z}_{\theta}^{f}+\mathbf{v};\theta)$ be the loss of generating $y^{f}$ given $\mathbf{z}_{\theta}^{f}+\mathbf{v}$ in model $f_{\theta}$. For simplicity, we write $\ell(y^{f}|\mathbf{z}_{\theta}^{f}+\mathbf{v})$ to present $\ell(y^{f}|\mathbf{z}_{\theta}^{f}+\mathbf{v};\theta)$. Following Koh and Liang [27], we assume that the loss is twice-differentiable and locally convex. Since $\mathbf{v}$ is small, we approximate the function $\ell(y^{f}|\mathbf{z}_{\theta}^{f}+\mathbf{v})$ using the second-order Taylor approximation:
$\ell(y^{f}|\mathbf{z}_{\theta}^{f}+\mathbf{v})\approx\ell(y^{f}|\mathbf{z}_{\theta}^{f})+\mathbf{v}^{\top}\nabla_{\mathbf{z}_{\theta}^{f}}\ell(y^{f}|\mathbf{z}^{f}_{\theta})+\frac{1}{2}\mathbf{v}^{\top}\nabla^{2}_{\mathbf{z}_{\theta}^{f}}\ell(y^{f}|\mathbf{z}_{\theta}^{f})\mathbf{v}$ (5)

### v4

@@ Abstract
Here, we show that current state-of-the-art LLM unlearning methods inherently reduce models’ robustness, causing them to misbehave even when a single non-adversarial forget-token is present in the retain-query. Toward understanding underlying causes, we propose a novel theoretical framework that reframes the unlearning process as backdoor attacks and defenses: forget-tokens act as backdoor triggers that, when activated in retain-queries, cause disruptions in unlearned models’ behaviors, similar to successful backdoor attacks. The sense that, LLM unlearning methods themselves poison the model, make it more vulnerable to forget-tokens, and hide rather than erase target knowledge, describes their true mechanism. To mitigate the vulnerability caused by the forgetting process, we reinterpret the retaining process as a backdoor defense and propose Random Noise Augmentation (RNA), a lightweight, model and method-agnostic approach with theoretical guarantees for improving the robustness of models. Extensive experiments demonstrate that RNA significantly improves the robustness of unlearned models while preserving forget and retain performances. This backdoor attack-defense framework offers insights into the mechanism of unlearning that can shed light on future research directions for improving unlearning robustness.
@@ 1 Introduction
Modern LLMs are pre-trained on massive text corpora and then post-trained with Reinforcement Learning from Human Feedback (Christiano et al., 2017; Ziegler et al., 2019; Stiennon et al., 2020; Ouyang et al., 2022) to be helpful and harmless (Bai et al., 2022). Recent studies have shown that despite safety enhancements, aligned LLMs can still exhibit harmful and undesirable behaviors, such as generating toxic content (Wen et al., 2023), producing copyrighted material (Karamolegkou et al., 2023; Eldan & Russinovich, 2023; Wei et al., 2024; Cooper et al., 2025), bias (Belrose et al., 2024), leaking sensitive and private information (Nasr et al., 2025; Patil et al., 2024), and potentially aiding malicious uses such as cyberattacks, chemical attacks, and bioweapons development (Fang et al., 2024; Sandbrink, 2023; Li et al., 2024). As LLMs advance in size and capabilities at an unprecedented speed, concerns about their potential risks continue to grow.
Machine Unlearning (MU; Cao & Yang (2015); Bourtoule et al. (2021); Nguyen et al. (2022); Xu et al. (2023)) is an approach aiming to robustly (1) remove specific target knowledge in a forget-set and capabilities from a pre-trained model, while (2) retaining the model’s other knowledge in a retain-set and capabilities. Recent works on the robustness of unlearning methods primarily focus on the first criterion, evaluating the robustness of unlearned models against knowledge recovery that adversarially tries to recover unlearned knowledge. For example, previously unlearned knowledge is shown to resurface through relearning (Li et al., 2024; Deeb & Roger, 2024; Lo et al., 2024), sequential unlearning (Shi et al., 2025), target relearning attacks (Hu et al., 2025), removing or steering specific directions in the latent space (Łucki et al., 2025; Seyitoğlu et al., ), quantization (Zhang et al., 2025), or even simply fine-tuning on unrelated tasks (Doshi & Stickland, 2024; Łucki et al., 2025).
However, the equally important criterion of robustly preserving the model’s general knowledge—that is, ensuring stable and accurate responses to retain-queries even when they inadvertently include forget-tokens—remains underexplored. Initial steps have been taken, such as Thaker et al. (2025), who examined the robustness of Representation Misdirection for Unlearning (RMU; Li et al. (2024)), demonstrating that RMU-unlearned models are fragile when asked with retain-queries (e.g., Q&A about general knowledge) containing forget-tokens (tokens in the forget-set). However, many critical questions remain unanswered. In this paper, we make the following contributions:
@@
➁We propose a novel perspective that decomposes the unlearning process into “forgetting” and “retaining” processes and reframes it as a backdoor attack and defense problem. The “forgetting” corresponds to a backdoor attack: by treating the forget-set as a poisoned dataset, we formulate how the unlearning methods inadvertently learn to align forget-tokens (backdoor triggers) with the target representations (the target labels). As a result, when forget-tokens appear in a retain-query, it is similar to activating the backdoor trigger, making the model misbehave. To counteract the vulnerability introduced by the “forgetting”, we reinterpret the “retaining” as a backdoor defense and introduce Random Noise Augmentation (RNA), a lightweight, model and method-agnostic approach which adds small, independent Gaussian noise to each retain-query’s representation during fine-tuning to reduce the model’s sensitivity to forget-tokens.
➂Through theoretical and empirical analysis, we show that RNA significantly improves the robustness of unlearned models while maintaining original forget and retain performances.
@@ 2 Related Works and Preliminaries
Machine Unlearning. MU has become one of the most important tools for ensuring the safety and protecting the privacy of LLMs (Xu et al., 2023; Nguyen et al., 2022; Barez et al., 2025). Most recent works on MU focus on developing algorithms for different tasks, domains, and settings (Pawelczyk et al., 2024; Thaker et al., 2024; Jin et al., 2024; Shi et al., 2025; Choi et al., 2024; Pal et al., 2025; Muhamed et al., 2025; Wang et al., 2025b; Kuo et al., 2025; Zhuang et al., 2025), while much less effort was spent on developing robust unlearning algorithms.
@@
Unlearning Robustness. Previous works on MU robustness focus on “forget-robustness”, studying the robustness of MU algorithms in making the model forget the target knowledge and capabilities. Researchers showed that unlearned knowledge can resurface through re-learning (Li et al., 2024; Lynch et al., 2024; Barez et al., 2025; Lo et al., 2024), sequential unlearning (Shi et al., 2025), quantization (Zhang et al., 2025), fine-tuning unlearned models on unrelated tasks (Doshi & Stickland, 2024; Łucki et al., 2025), and adversarial attacks (Hu et al., 2025; Yuan et al., 2025a; Shumailov et al., 2024; Huang et al., 2025; Wu et al., 2025) and developed methods for improving forget-robustness of MU algorithms (Sheshadri et al., 2024; Tamirisa et al., 2025; 2024; Fan et al., 2025; Zhang et al., 2025; Wang et al., 2025a). This paper explores the “retain-robustness” of MU algorithms, studying the robustness of MU algorithms in robustly retaining the original model’s general knowledge and capabilities. Thaker et al. (2025) presented preliminary results showing that state-of-the-art MU algorithms do not preserve the original model’s knowledge and capabilities. We bridge the gap in retain-robustness research by introducing RNA, a simple data augmentation method inspired by adversarial training to improve the robustness of MU algorithms.
@@
Notation and problem formulation. The training data of an MU problem consists of two subsets: the forget-set $\mathcal{D}_{f}$ and the retain-set $\mathcal{D}_{r}$. The goal is to minimize the model’s performance on the forget set while keeping the performance on the retain-set. Let $f_{\bm{\theta}}$ be a model parameterized by $\bm{\theta}$, and $\ell(\mathbf{y}|\mathbf{x};\bm{\theta})$ is the loss of input $\mathbf{x}$ with respect to a target output $\mathbf{y}$ in model $f_{\bm{\theta}}$. A commonly used form of unlearning involves minimizing the following two-part loss:
$\mathcal{L}_{\mathcal{D}_{f},\mathcal{D}_{r},\bm{\theta}}=\mathbb{E}_{(\mathbf{x}^{f},\mathbf{y}^{f})\sim\mathcal{D}_{f}}\left[\ell\left(\mathbf{y}^{f}|\mathbf{x}^{f};\bm{\theta}\right)\right]+\alpha\mathbb{E}_{(\mathbf{x}^{r},\mathbf{y}^{r})\sim\mathcal{D}_{r}}\left[\ell\left(\mathbf{y}^{r}|\mathbf{x}^{r};\bm{\theta}\right)\right]$ (1)
@@
where $\mathbf{y}^{f},\mathbf{y}^{r}$ are the target outputs of forget and retain input, respectively, $\alpha\in\mathbb{R}^{+}$ is a retain weight. We consider two widely used classes of LLM unlearning methods, which rely on Representation Misdirection (RM) and Preference Optimization (PO). We denote $||\cdot||$ the Euclidean norm.
2.1 Representation Misdirection
@@
Representation Misdirection (RMU and its variants) is an unlearning approach that conducts unlearning by manipulating latent representations during fine-tuning. Denote $\mathbf{z}_{\bm{\theta}}^{f},\mathbf{z}_{\bm{\theta}}^{r}\in\mathbb{R}^{n\times d_{l}}$ the output hidden state of $n$-tokens in forget-sample $\mathbf{x}^{f}$ and in retain-sample $\mathbf{x}^{r}$, respectively, at layer $l$ in model $f_{\bm{\theta}}$, where $d_{l}$ is the dimension of hidden states at layer $l$. In this context, the activations from the MLP module in the transformer layer are used as the latent representations.
@@
Representation Misdirection for Unlearning (RMU; Li et al. (2024)) pushes the latent representation of forget-tokens to a predetermined random representation $\mathbf{y}^{f}=c\mathbf{u}$, where $\mathbf{u}\in\mathbb{R}^{d_{l}}$ is a unit vector with each element uniformly sampled from $[0,1)$, and $c\in\mathbb{R}^{+}$ is a coefficient. It also regularizes the latent representation of retain-tokens back to the reference model’s representation:
[…]
@@
where $\bm{\theta}$ and $\bm{\theta}^{\text{ref}}$ are the parameters of the updated and reference (frozen weight) models, respectively.
[…]
@@
Random Steering Vector (RSV). Additionally, we implement RSV—a variant of RMU that uses the target random representation $\mathbf{y}^{f}=\mathbf{z}^{f}_{\theta^{\text{ref}}}+c\bm{\epsilon}$, where $c\in\mathbb{R}^{+}$ is a predetermined coefficient, $\bm{\epsilon}$ is a random unit vector sampled from Gaussian distribution $\mathcal{N}(\bm{0},\mu\bm{I})$, $\mu\bm{I}$ is covariance matrix, $\mu\in\mathbb{R}^{+}$.
2.2 Preference Optimization
@@
Negative Preference Optimization (NPO; Zhang et al. (2024)) treats forget-samples as negative preference samples in Direct Preference Optimization (DPO; Rafailov et al. (2023)). NPO can be viewed as a gradient ascent variant with adaptive gradient weights that allows more controlled and stable optimization:
@@
$\mathcal{L}^{\text{NPO}}=\mathbb{E}_{(\mathbf{x}^{f},\mathbf{y}^{f})\sim\mathcal{D}_{f}}\left[-\frac{2}{\beta}\log\sigma\left(-\beta\log\left(\frac{\pi_{\bm{\theta}}(\mathbf{y}^{f}|\mathbf{x}^{f})}{\pi_{\bm{\theta}^{\text{ref}}}(\mathbf{y}^{f}|\mathbf{x}^{f})}\right)\right)\right],$ (3)
@@
where $\beta\in\mathbb{R}^{+}$ is a temperature hyperparameter (NPO reduces to gradient ascent as $\beta\to 0$), $\sigma(\cdot)$ is the sigmoid function, and $\pi_{\bm{\theta}}(\mathbf{y}^{f}|\mathbf{x}^{f})$, $\pi_{\bm{\theta}^{\text{ref}}}(\mathbf{y}^{f}|\mathbf{x}^{f})$ denotes the predicted probability of $\mathbf{y}^{f}$ given $\mathbf{x}^{f}$ in the model $f_{\bm{\theta}}$ and reference model $f_{\bm{\theta}^{\text{ref}}}$ (frozen weight) respectively.
@@
Simple Negative Preference Optimization (SimNPO; Fan et al. (2024)) simplifies NPO by using a normalized sequence log-probability and introducing a reward margin hyperparameter $\gamma\geq 0$:
@@
$\mathcal{L}^{\text{SimNPO}}=\mathbb{E}_{(\mathbf{x}^{f},\mathbf{y}^{f})\sim\mathcal{D}_{f}}\left[-\frac{2}{\beta}\log\sigma\left(-\frac{\beta}{|\mathbf{y}^{f}|}\log\pi_{\theta}(\mathbf{y}^{f}|\mathbf{x}^{f})-\gamma\right)\right],$ (4)
where $|\mathbf{y}^{f}|$ is the length of output $\mathbf{y}^{f}$.
@@
Direct Preference Optimization (DPO). As a baseline, Zhang et al. (2024); Maini et al. (2024); Yuan et al. (2025b)) adopted standard DPO, using a refusal answer $\mathbf{y}^{\text{idk}}\in\mathcal{D}_{\text{idk}}$ such as “I Don’t Know” as the positive samples and forget-samples as negative samples.
[…]
@@ 3 A Unified View of LLM Unlearning
We first draw a connection between RM and PO methods through a unified view of the generative latent variable model (GLVM). Let $\mathbf{z}_{\bm{\theta}}^{f}+\mathbf{v}$ be the steered latent representation of forget-sample $\mathbf{x}^{f}$ in $f_{\bm{\theta}}$ as a result of RM. We assume that random vector $\mathbf{v}$ is small and sampled from normal distribution $\mathcal{N}(\bm{0},\mu\bm{I})$, $\mu\in\mathbb{R}^{+}$. We employ the notion of the GLVM, that is, GLVM $f_{\bm{\theta}}$ generates target output $\mathbf{y}^{f}$ given the latent variable $\mathbf{z}_{\bm{\theta}}^{f}$. Let $\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f}+\mathbf{v};\bm{\theta})$ be the loss of generating $\mathbf{y}^{f}$ given $\mathbf{z}_{\bm{\theta}}^{f}+\mathbf{v}$ in model $f_{\bm{\theta}}$. For simplicity, we write $\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f}+\mathbf{v})$ to present $\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f}+\mathbf{v};\bm{\theta})$. Following Koh & Liang (2017), we assume that the loss is twice-differentiable and locally convex. Since $\mathbf{v}$ is small, we approximate the function $\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f}+\mathbf{v})$ using the second-order Taylor approximation:
$\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f}+\mathbf{v})\approx\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f})+\mathbf{v}^{\top}\nabla_{\mathbf{z}_{\bm{\theta}}^{f}}\ell(\mathbf{y}^{f}|\mathbf{z}^{f}_{\bm{\theta}})+\frac{1}{2}\mathbf{v}^{\top}\nabla^{2}_{\mathbf{z}_{\bm{\theta}}^{f}}\ell(\mathbf{y}^{f}|\mathbf{z}_{\bm{\theta}}^{f})\mathbf{v}$ (5)